Calculate the concentration of buffer components present in 177.00 mL of a buffer solution that contains  0.350 M NH4Cl and 0.350 M NH3 immediately after the addition of 1.30 mL of 6.00 M HNO3.

Part 1:

[NH3]=

Part 2:

[NH4+]
Answer #1

mol of HNO3 added = 6.0M *1.3 mL = 7.8 mmol

NH3 will react with H+ to form NH4+

Before Reaction:

mol of NH3 = 0.35 M *177.0 mL

mol of NH3 = 61.95 mmol

mol of NH4+ = 0.35 M *177.0 mL

mol of NH4+ = 61.95 mmol

after reaction,

mol of NH3 = mol present initially - mol added

mmol of NH3 = (61.95 - 7.8) mmol

mol of NH3 = 54.15 mmol

mol of NH4+ = mol present initially + mol added

mol of NH4+ = (61.95 + 7.8) mmol

mol of NH4+ = 69.75 mmol

Total volume = 177 mL + 1.30 mL = 178.30 mL

So,

[NH3] = mol of NH3 / total volume

= 54.15 mmol / 178.30 mL

= 0.304 M

[NH4+] = mol of NH4+ / total volume

= 69.75 mmol / 178.30 mL

= 0.391 M

Answer:

[NH3] = 0.304 M

[NH4+] = 0.391 M